OpenStudy (anonymous):

plz help me with this one, plz provide the work or start with taking the derivative/double derivative of this. y=x^2 +1/x^2 -1 find the relative extrema, points of inflection, intervals where increasing/decreasing, intervals where concave up/down and any vertical or horizontal asym. show all work and sign charts, and tests used.

OpenStudy (anonymous):

@DanJS

zepdrix (zepdrix):

Hey Katie :) The problem looks like this? \[\large\rm y=\frac{x^2+1}{x^2-1}\]Quotient rule would be a good place to start, ya?

OpenStudy (anonymous):

hi yes, looks just like that and yeah use quotient rule for that and take the first derivitive to start with

zepdrix (zepdrix):

I like to "set up" the rule before applying it. Nice way to stay organized, at least early on in calculus.\[\large\rm y'=\frac{\color{royalblue}{(x^2+1)'}(x^2-1)-(x^2+1)\color{royalblue}{(x^2-1)'}}{(x^2-1)}\]So the blue stuff is what we need to differentiate, ya? Simple power rule gives us:\[\large\rm y'=\frac{\color{orangered}{(2x)}(x^2-1)-(x^2+1)\color{orangered}{(2x)}}{(x^2-1)}\]

OpenStudy (anonymous):

yes, i got that part. neat work i must say!

zepdrix (zepdrix):

Woops! Denominator should've been squared :) silly mistake on my part.

OpenStudy (anonymous):

yes (x^2 -1)^2 sorry i should have pointed it out

zepdrix (zepdrix):

For extrema, we set our first derivative to zero,\[\large\rm 0=\frac{(2x)(x^2-1)-(x^2+1)(2x)}{(x^2-1)^2}\] The numerator is going to give us our traditional `critical points`, locations of horizontal tangency, while the denominator will tell us where the function has vertical tangency. \(\large\rm 0=(x^2-1)^2\qquad\to\qquad x=\pm1\) Notice that x=1 and x=-1 are not in the `domain` of this function, so we ignore those values here.

zepdrix (zepdrix):

Hmm they wanted us to talk about asymptotes? Maybe we should have done that first, before taking derivatives. Either way is fine I suppose :d

zepdrix (zepdrix):

When we set our numerator equal to zero,\[\large\rm 0=(2x)(x^2-1)-(x^2+1)(2x)\]What critical points do you get? :)

OpenStudy (anonymous):

okay wait let me tell u the HA and VA that i got

zepdrix (zepdrix):

Oh sure.

OpenStudy (anonymous):

y=1 HA?

zepdrix (zepdrix):

\[\large\rm \lim_{x\to\infty}y=1\]Ok that sounds right so far! As we go far far to the right (or left), our function is approaching this line y=1.

OpenStudy (anonymous):

okay and VA in this case would be 0+1/0-1 so 0?

OpenStudy (anonymous):

wait sorry that's the y-inter

OpenStudy (anonymous):

okay VA= 1 and -1

zepdrix (zepdrix):

Denominator can never be zero,\[\large\rm x^2-1\ne0\qquad\implies\qquad (x-1)(x+1)\ne0\]\[\large\rm x\ne1,\qquad x\ne-1\]Ok great.

OpenStudy (anonymous):

okay lets cont with the direvitives! tysm so far

zepdrix (zepdrix):

You have your first derivative, setting it equal to zero so we can find critical points,\[\large\rm 0=\frac{(2x)(x^2-1)-(x^2+1)(2x)}{(x^2-1)^2}\]Multiply both sides by (x^2-1)^2 to get rid of the denominator,\[\large\rm 0=(2x)(x^2-1)-(x^2+1)(2x)\]Now solve for x, what do you get for critical points? :)

OpenStudy (anonymous):

(2x^3 -2x)- (2x^3+2x)

zepdrix (zepdrix):

Hmmm, you can do that I suppose. But I would instead recommend factoring, unless that's going to be too confusing.

zepdrix (zepdrix):

\[\large\rm 0=\color{royalblue}{(2x)}(x^2-1)-(x^2+1)\color{royalblue}{(2x)}\]\[\large\rm 0=\color{royalblue}{(2x)}\left[(x^2-1)-(x^2+1)\right]\]

zepdrix (zepdrix):

And then apply your `Zero-Factor Property`:\[\large\rm 0=2x,\qquad\qquad 0=(x^2-1)-(x^2+1)\]

OpenStudy (anonymous):

x=0 and x=1 and x=-1

zepdrix (zepdrix):

x=0 sounds good! 1 and -1, hmmm let's see... distributing that negative sign gives us:\[\large\rm 0=x^2-1-x^2-1\]The x^2's subtract out,\[\large\rm 0=-2\]

zepdrix (zepdrix):

So it looks like we're not getting any critical points from this factor, ya? :o

OpenStudy (anonymous):

wait i am reading over your stuff

OpenStudy (anonymous):

no, the ones factored are our zeros right

OpenStudy (anonymous):

cause numberator are the critical zeros right

zepdrix (zepdrix):

Zero-Factor Property tells us when we have things multiplied together giving us zero, we can set each individual factor equal to zero:\[\large\rm 0=\color{royalblue}{(2x)}\color{orangered}{\left[(x^2-1)-(x^2+1)\right]}\] \[\large\rm \color{royalblue}{0=2x},\qquad\qquad \color{orangered}{0=\left[(x^2-1)-(x^2+1)\right]}\]This blue piece gave you x=0, good. I'm trying to say that this orange piece will not give you x=1 and -1. Hopefully the algebra I did before was clear, I skipped a couple steps maybe.

zepdrix (zepdrix):

The orange is actually giving us `no critical points`. We only have the critical point at x=0.

zepdrix (zepdrix):

Let's see if we can figure out increasing/decreasing :)

zepdrix (zepdrix):

brb one sec, i need some peanut butter :O

OpenStudy (anonymous):

right i see what u are saying. x^2 -1 -x^2-1=0 the X^2 got cancled but the -1 and -1 becomes -2 so because the x are alredy cancled right?

OpenStudy (anonymous):

yeah go ahead i am not leaving untill i get this solved. haha

OpenStudy (anonymous):

tysm for helping me! :)

zepdrix (zepdrix):

If you're more familiar with doing a sign chart, we can try that. I don't quite remember how to do that though... I always did the interval method. They're essentially the same though, I think.

OpenStudy (anonymous):

yes sign charts plz i am familiar

zepdrix (zepdrix):

\[\large\rm y'=\frac{(2x)(x^2-1)-(x^2+1)(2x)}{(x^2-1)^2}\]Notice that our denominator is being squared. It will ALWAYS be positive. That will have no effect on the `sign` of our derivative, so we can ignore it.\[\large\rm y'=(2x)(x^2-1)-(x^2+1)(2x)\]This is kind of sloppy notation saying that y' equals this ^ But whatever, it simplifies things for us.

zepdrix (zepdrix):

We need to know what is happening on the left and right side of .... zero, since that is our critical point.

OpenStudy (anonymous):

right but did u read my commet about -2 is that right the way i explained that x^2 gets cancled that's why we do not get a critical point there?

zepdrix (zepdrix):

Hey @blueskies :) Don't hijack someone else's question. I see you have your question open. In the chat box below your question type @zepdrix or someone else's name like @Jhannybean and it will alert them that you need assistance. I'll try to take a look if I get a chance.

OpenStudy (anonymous):

thank you @zepdrix :)

zepdrix (zepdrix):

Ya that's a good way to explain it Katie. You actually end up with 0=-2 after you cancel out the x^2's. And this is a FALSE statement. 0 can not equal -2, so this is giving us no information, we can't have a zero here.

OpenStudy (anonymous):

right so that concludes 0 being the only critical point. okay you can carry the sign chart now. sorry to ask that agian.

zepdrix (zepdrix):